A rectangle has a length of 38 yards less than four times it's width. The area of the rectangle is 840 yd.² find the length of the rectangle
Answer by jorel1380(3719)

Let w be the width. Then the length would be 4w-38. So:
w(4w-38)=840
4w²-38w-840=0
2w²-19w-420=0
(2w+21)(w-20)=0
w=-21/2 or 20
The width is 20; the length is 42
☺☺☺☺